OpenStudy (anonymous):

functions: If f(x)=x^2-x+1, solve: 1) f(x+1) 2) f(-a) 3) 2f(x)-2 4) f(2x)

OpenStudy (anonymous):

CAN I post my answers and someone check it?

OpenStudy (anonymous):

you mean "evaluate"?

OpenStudy (anonymous):

probably, sorry

OpenStudy (anonymous):

yeah, lol, just substitute whatever is in f(whatever is in here)

OpenStudy (anonymous):

yes i'll check it

OpenStudy (anonymous):

1) x^2+x+3 2) -a^2+a+1 (i didn't know if there was any way to simplify this? 3) 2x^2-x 4) 2x^2-2x+1

OpenStudy (anonymous):

all wrong, it three right 2 f(x)-2?

OpenStudy (anonymous):

its 2*f(x) -2

OpenStudy (anonymous):

can you show me how to do them then?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

jaersyn i got this

OpenStudy (anonymous):

all yours kmalone99

OpenStudy (anonymous):

<_<

OpenStudy (anonymous):

\[1) f(x+1)= (x+1)^{2}-(x+1)+1\] \[=(x+1)(x+1)-x-1+1\] \[=x ^{2}+2x+1-x=x ^{2}+x+1\]

OpenStudy (anonymous):

Ok, I get #1. I just didn't distribute the negative sign.

OpenStudy (anonymous):

2) f(-a)=\[f(-a)=(-a)^{2}-(-a)+1=a ^{2}+a+1\]

OpenStudy (anonymous):

now retry 3 and 4

OpenStudy (anonymous):

i got 2... i dont see what went wrong with 3 and i dont know how to simplify 4

OpenStudy (anonymous):

wait- 3 should be 2x^2-2x?

OpenStudy (anonymous):

great now on four you just did not square the 2

OpenStudy (anonymous):

oh! so it would be 4x^2-2x+1 ?

OpenStudy (anonymous):

yes great job you get A+

OpenStudy (anonymous):

yay! lol I just always make silly simplifying mistakes or I miss a sign or something simple like that. Oh well, thanks very much ^_^